a kite is flying on 50ft of string. its vertical distance form the ground is 10ft more than its horizontal distance form the person flying it. assuming that the string is being held at ground level, find its horizontal distance from the person and its vertical distance from the ground.
Answer by ankor@dixie-net.com(22740)

a kite is flying on 50ft of string. its vertical distance form the ground is 10ft more than its horizontal distance form the person flying it.
assuming that the string is being held at ground level, find its horizontal distance from the person and its vertical distance from the ground.
:
This is a pythag problem a^2 + b^2 = c^2
let x = horizontal distance
then
(x+10) = the vertical distance
so we have
x^2 + (x+10)^2 = 50^2
x^2 + x^2 + 20x + 100 = 2500
2x^2 + 20x + 100 - 2500 = 0
2x^2 + 20x - 2400 = 0
Simplify, divide by 2
x^2 + 10x - 1200 = 0
Factors to
(x+40)(x-30) = 0
the positive solution
x = 30 ft is the horizontal distance
You can find the vertical distance